# Enhancing Data Lakes with GraphAr: Efficient Graph Data Management with a Specialized Storage Scheme

### v3

@@ 6.4 Storage Media
We assess the efficiency of GraphAr across various storage media: local in-memory tmpfs, ESSD (an Alibaba Cloud virtualized elastic block device), and S3-like Object Store Service (OSS). The graph used is SF100, with specifically foucus on the comment vertex type and comment_hasTag_tag edge type.
Table 2 encapsulates the efficiency of GraphAr across different storage media. These results conclusively establish that GraphAr is not only efficient but also robust, delivering consistently high performance, with speedups of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
6.5 End-to-end Workloads
To demonstrate the practicality of GraphAr in real-world scenarios, we conduct a performance evaluation using end-to-end workloads from the LDBC SNB benchmark [32, 57], which is widely used in the graph processing community. Although the benchmark specifies vertex/edge types, it does not explicitly define the labels. However, we are able to identify certain vertex types that are static (e.g., tagclass and place), which have a fixed and very small vertex set size that does not scale with the graph size. On the other hand, vertex types like comment and person are considered dynamic. Based on this observation, we can treat information related to static types as labels for dynamic types in GraphAr, for example, all tag classes of a comment are attached as labels for the corresponding comment vertex. Similar strategies are also adopted by graph databases [37] to optimize data access performance.
For evaluation, graphs at different scales (listed in Table 1) are generated using the LDBC SNB data generator, as Parquet files. These graphs are then converted into GraphAr format, with the vertex labels attached as described above. Upon investigating the benchmark, including $7$ short and $14$ complex interactive queries, as well as $20$ business intelligence queries, we find that neighbor retrieval is frequently encountered, involved in approximately $90\%$ of the queries. Considering the aforementioned label organization, label filtering is also common, involved in approximately $50\%$ of the queries.
Table 2: Performance comparison across storage media.
@@ 6.5 End-to-end Workloads
6.5.1 Query Implementations
@@
Queries. The evaluation focuses on three representative queries, with the required parameters set according to the reference implementations [12, 11]. The first query, IS-3 (interactive-short-3), aims to find all the friends of a given person and return their information. It exemplifies the common pattern of querying neighboring vertices and retrieving associated properties. The second query, IC-8 (interactive-complex-8), is more complex as it involves traversing multiple hops from the starting vertex. Lastly, the BI-2 (business-intelligence-2) query aims to analyze tag evolution. It involves finding and counting the messages associated with tags within a specific tag class, thus requiring vertex filtering by labels.
@@
GraphAr. We develop hand-written implementations for each query based on GraphAr, which utilize the data organization and specifically prioritize two essential operations: neighbor retrieval and label filtering. Our implementation refers to the official reference implementations [12, 11] in Cypher or SQL, to ensure equivalence to the original queries.
@@
Acero. We also implement these queries in Acero [1], which is a powerful C++ library integrated into Apache Arrow for analyzing large streams of data. It offers a comprehensive set of operators such as scan, filter, project, aggregate, and join, among others. Moreover, Acero supports taking Parquet as the data source and enables the pushdown of predicates, making it a strong baseline for comparison with GraphAr. Despite our best efforts to optimize it, we do not perform data re-organization or utilize GraphAr’s encoding/decoding optimizations for the implementations based on Acero.
6.5.2 Performance Comparison
@@
Table 3 presents a comparison of end-to-end time performance, demonstrating that the implementation based on GraphAr significantly outperforms Acero, achieving an average speedup of 29.5×29.5$\times$29.5 ×. Upon closer examination of the results, it becomes apparent that the performance improvement can be attributed to the following factors: 1) data layout design and encoding/decoding optimizations we proposed, to enable efficient neighbor retrieval (IS-3, IC-8, BI-2) and label filtering (BI-2), as demonstrated in micro-benchmarks; 2) bitmap generation during the two critical operations, which can be utilized in subsequent selection steps (IS-3, IC-8, BI-2).
Taking IS-3 as an example, the optimization of neighbor retrieval achieves an average speedup of 571.4×571.4$\times$571.4 × compared to the baseline utilizing Parquet’s filter pushdown. The subsequent selection based on bitmaps provides a speedup of 2.7×2.7$\times$2.7 ×. Other steps like order_by show similar performance in GraphAr and Acero due to closely matched implementations.
Table 3: Execution time for IS-3, IC-8 and BI-2 (in seconds).
Graph SF30 SF100 SF300
Impl. Acero GraphAr Acero GraphAr Acero GraphAr
IS-3 0.156 0.005 0.475 0.010 1.390 0.029
IC-8 72.22 3.362 245.5 6.563 894.4 23.29
BI-2 67.74 4.295 231.6 16.28 755.6 50.04
@@
  • Storage efficiency: GraphAr remarkably reduces storage requirements, using only $27.3\%$ of the storage compared to baseline methods for storing topology, and as low as $2.9\%$ for label storage on average.
  • Query performance: GraphAr significantly outpaces the baselines in retrieval time, achieving speedups ranging from 49.8×49.8$\times$49.8 × to 9232.4×9232.4$\times$9232.4 × for neighbor retrieval and an average speedup of 6.0×6.0$\times$6.0 × for simple label filtering, as observed in micro-benchmarks on the ESSD storage.
  • Storage media: Preliminary evaluations indicate seamless compatibility across various storage layers like local in-memory tmpfs, ESSD, and S3-like Object Store Service (oss), all achieving high speedup of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
@@
7 Related Work
File formats in data lakes. The data lake ecosystem encompasses various common file formats, including CSV, JSON, Protocol Buffers, HDF5 [18], AVRO [3], ORC [4], and Parquet [5]. While these formats support various optimizations that benefit both tables and graphs, they do not inherently cater to some unique needs of LPGs, thus fall short in representing LPG semantics and supporting graph-specific operations.
Data management in data lakes. The popularity of data lakes has led to efforts aimed at enhancing their architecture and data management [48, 25, 16, 54, 49]. These endeavors primarily focus on managing existing data within data lakes and are distinct from GraphAr. GraphAr, on the other hand, can be considered as a new storage format with unique features. It can be leveraged by these existing works to further extend the utility and capabilities of data lakes.
Graph file formats. There are specific formats designed for graph data [27, 38, 9, 24] and RDF (Resource Description Framework) data [17, 46]. However, the primary focus of these formats is to describe or exchange graph data in a standardized manner, e.g., utilizing XML, and are not optimized for storage and retrieval purposes. The lack of encoding, compression and push-down optimizations can lead to far inferior performance, making them less suitable for managing LPGs in data lakes.
Graph databases. There are also graph databases [13, 31, 10] that are designed to store and manage graph data. While they offer various features that are tailored for LPGs, they primarily focus on in-memory mutable data management, operating at a higher level compared to GraphAr. GraphAr, with its format compatible with the LPG model, can be utilized as an archival format for graph databases.
Operation pushdown. Some previous works [45, 60, 61, 62, 63, 6] aim to develop high-performance operators on storage formats of either column-oriented or row-oriented. These works and GraphAr share the same goal of improving pushdown operators and making the utilization of available CPU instructions. However, it is important to note that these works focus on operations related to relational data, such as scan, select, and filter based on properties. In contrast, GraphAr specifically focuses on two graph-specific operations.
8 Conclusion

### v4

@@ 6.4. Storage Media
We assess the efficiency of GraphAr across various storage media: local in-memory tmpfs, ESSD (an Alibaba Cloud virtualized elastic block device), and S3-like Object Store Service (OSS). The graph used is SF100, with specifically focus on the comment vertex type and comment_hasTag_tag edge type. Table 2 encapsulates the efficiency of GraphAr across different storage media. These results demonstrates that GraphAr is not only efficient but also robust, delivering consistently high performance, with speedups of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
Table 2. Performance comparison across storage media.
@@
6.5. End-to-end Graph Query Workloads
To demonstrate the practicality of GraphAr in real-world scenarios, we conduct a performance evaluation using end-to-end workloads from the LDBC SNB benchmark (Erling et al., 2015; Szárnyas et al., 2022), which is widely used in the graph processing community. Although the benchmark specifies vertex/edge types, it does not explicitly define the labels. However, we are able to identify certain vertex types that are static (e.g., tagclass and place), which have a fixed and very small vertex set size that does not scale with the graph size. On the other hand, vertex types like comment and person are considered dynamic. Based on this observation, we can treat information related to static types as labels for dynamic types in GraphAr, for example, all tag classes of a comment are attached as labels for the corresponding comment vertex. Similar strategies are also adopted by graph databases (Gupta et al., 2021) to optimize data access performance.
For evaluation, graphs at different scales (listed in Table 1) are generated using the LDBC SNB data generator. These graphs are then converted into GraphAr format, with the vertex labels attached as described above. Upon investigating the benchmark, including $7$ short and $14$ complex interactive queries, as well as $20$ business intelligence queries, we find that neighbor retrieval is frequently encountered, involved in approximately $90\%$ of the queries. Considering the aforementioned label organization, label filtering is also common, involved in approximately $50\%$ queries.
@@ 6.5. End-to-end Graph Query Workloads
Query implementations. The evaluation focuses on three representative queries, with the required parameters set according to the reference implementations (ldb, 2024; bi_, 2024). IS-3 (interactive-short-3) aims to find all the friends of a given person and return their information. It exemplifies the common pattern of querying neighboring vertices and retrieving associated properties. IC-8 (interactive-complex-8) is more complex as it involves traversing multiple hops from the starting vertex. Lastly, the BI-2 (business-intelligence-2) query involves finding and counting the messages associated with tags within a specific tag class, thus requiring vertex filtering by labels.
@@
We develop hand-written implementations for each query based on GraphAr, which utilize the data organization and specifically prioritize two essential operations: neighbor retrieval and label filtering. Our implementation adheres to the official reference implementations (ldb, 2024; bi_, 2024) to ensure equivalence to the original queries.
@@
We then implement these queries in Acero (Ace, 2024), which is a powerful C++ library integrated into Apache Arrow for analyzing large streams of data. It offers a comprehensive set of operators such as scan, filter, project, aggregate, and join, among others. Moreover, Acero supports taking Parquet as the data source and enables the pushdown of predicates, making it a strong baseline for comparison with GraphAr. Despite our best efforts to optimize it, we do not perform data re-organization or utilize GraphAr’s encoding/decoding optimizations for this implementation based on Acero.
We also include two additional baselines: Apache Pinot (pin, 2024), a real-time OLAP datastore used by LinkedIn for processing and querying large social networks, and Neo4j (neo, 2024b), a main graph database utilizing the Cypher query language. While both are widely-used, they are not natively designed for data lakes and require an Extract-Transform-Load (ETL) process for integration.
Table 3. Query execution times (in seconds), with the format of Pinot (P), Neo4j (N), Acero (A), GraphAr (G). “OM” denotes failed execution due to out-of-memory errors.
SF30 SF100 SF300
P N A G P N A G P N A G
ETL 6024 390 17726 2094 OM 9122
IS-3 1.00 0.30 0.16 0.01 6.59 2.09 0.48 0.01 OM 4.12 1.39 0.03
IC-8 1.35 0.37 72.2 3.36 8.43 1.26 246 6.56 OM 2.98 894 23.3
BI-2 125 45.0 67.7 4.30 3884 1101 232 16.3 OM 6636 756 50.0
@@
Performance comparison. Table 3 presents a comparison of end-to-end performance, clearly demonstrating that the implementation based on GraphAr significantly outperforms Acero, achieving an average speedup of 29.5×29.5$\times$29.5 ×. A closer analysis of the results reveals that the performance gains stem from the following factors: 1) data layout design and encoding/decoding optimizations we proposed, to enable efficient neighbor retrieval (IS-3, IC-8, BI-2) and label filtering (BI-2), as demonstrated in micro-benchmarks; 2) bitmap generation during the two critical operations, which can be utilized in subsequent selection steps (IS-3, IC-8, BI-2).
As for Pinot and Neo4j, their end-to-end performance is often dominated by extensive ETL processes, in the context of data lakes, as the results show. GraphAr performs best on IS-3, which is a single-hop query, and BI-2, where GraphAr utilizes label filtering for the early elimination of irrelevant data. While on IC-8, GraphAr is outperformed by Neo4j due to the query involving traversing multiple hops, which results in a significant volume of data loading for both Acero and GraphAr. Nevertheless, GraphAr not only offers efficient query performance but also eliminates the ETL overhead, potentially avoids out-of-memory errors that may occur. Thus, GraphAr provides a more practical solution for data lake scenarios.
6.6. Integration with Graph Processing Systems
One of the advantages of GraphAr is its compatibility with existing graph processing systems, enabling seamless integration. To demonstrate this, we have successfully integrated GraphAr into GraphScope (Fan et al., 2021) as the archive format for persistent storage. In addition, it is also utilized as an accessible storage backend for executing infrequent queries in an out-of-core manner. GraphScope is a distributed system for addressing a wide array of graph computing need, which has achieved significant technological advancements and gained widespread adoption across various industries.
Serve as the archive format. We first compare the performance of building graphs in GraphScope using 8 nodes, from external storages in GraphAr format against the baseline, where the datasets are in CSV format, sourced directly from the data providers. The findings illustrated in Figure 10(a) indicate that GraphAr significantly outperforms the baseline, achieving an average speedup of 4.9×4.9$\times$4.9 ×. This improvement can be attributed to GraphAr’s efficient encoding strategies that reduce the data volume to be loaded, as well as its optimized data organization and layout, which facilitate a faster in-memory graph construction within GraphScope.
Serve as a storage backend. Leveraging the capabilities for graph-related querying, the graph query engine within GraphScope can execute queries directly on the GraphAr data in an out-of-core manner. We evaluate the performance of GraphScope with GraphAr as the storage backend, and compare the average querying time of BI queries on SF30 with two baseline scenarios wherein GraphScope relies on its native in-memory storage options, specifically the immutable (“Imm”) and mutable (“Mut”) variants. Figure 10(b) demonstrates that although the querying time with GraphAr exceeds that of the in-memory storages, attributable to intrinsic I/O overhead, it significantly surpasses the process of loading and then executing the query, by 2.4×2.4$\times$2.4 × and 2.5×2.5$\times$2.5 ×, respectively. This indicates that GraphAr is a viable option for executing infrequent queries.
Application scenarios. This integration demonstrates the potential benefits of GraphAr in improving the efficiency of graph processing systems. In summary, its application scenarios include: 1) Data loading: GraphAr can significantly reduce graph loading times, making it an ideal choice for external storage formats; 2) Out-of-core queries: GraphAr can serve as a storage backend for executing graph queries in an out-of-core manner. It is particularly beneficial for infrequent queries that access only a portion of the graph data, eliminating the need for full in-memory graph representation. It also enables querying graphs that exceed the capacity of available memory. In other scenarios, such as: 1) real-time queries requiring low latency and frequent execution, or 2) graph analytics algorithms involving iterative computations across the entire graph (e.g., PageRank), GraphAr might not be the optimal direct storage solution and in-memory storage options are more suitable.
(a) Graph loading time.
(b) BI queries execution time.
Figure 10. GraphScope’s performance w/ and w/o GraphAr.
@@ 6.6. Integration with Graph Processing Systems
  • Storage efficiency: GraphAr remarkably reduces storage requirements, using only $29.2\%$ of the storage compared to baseline methods for storing topology, and as low as $2.5\%$ for label storage on average.
  • Query performance: GraphAr significantly outpaces the baselines in retrieval time, achieving an average speedup of 4452×4452$\times$4452 × for neighbor retrieval, and an average speedup of 14.8×14.8$\times$14.8 × for simple label filtering, as observed in micro-benchmarks on the ESSD storage.
  • Storage media: Evaluations indicate seamless compatibility across various storage layers like tmpfs, ESSD, and oss, all achieving high speedup of $\times8$ to 154×154$\times$154 × for neighbor retrieval and 2.7×2.7$\times$2.7 × to $\times1$ for label filtering.
@@
  • Compatibility: GraphAr is seamlessly integrated into a widely-used graph processing system GraphScope, enhancing its graph loading efficiency by 4.9×4.9$\times$4.9 ×, and accelerating its infrequent query execution with a speedup of 2.4×2.4$\times$2.4 ×.
@@
7. Related Work
File formats in data lakes. The data lake ecosystem encompasses various common file formats, including CSV, JSON, Protocol Buffers, HDF5 (hdf, 2024), AVRO (avr, 2024), ORC (orc, 2024), and Parquet (par, 2024). While these formats support various optimizations that benefit both tables and graphs, they fall short in comprehensively representing LPG semantics and supporting graph-specific operations.
Data management in data lakes. The popularity of data lakes has led to efforts aimed at enhancing their architecture and data management (Nargesian et al., 2019; Behm et al., 2022; pup, 2024; Ravat and Zhao, 2019; Neunhöffer, 2024). As for LPG management in data lakes, LinkedIn uses Apache Gobblin (gob, 2024) for data ingestion and employs Apache Spark (Zaharia et al., 2016) and Apache Pinot (pin, 2024) for processing large graph datasets representing the social network. Graph-specific querying frameworks like Neo4j (neo, 2024b) and Apache TinkerPop (tin, 2024) are also widely-used, and they integrate with data lakes via ETL processes. These endeavors primarily focus on managing existing data within data lakes and are distinct from GraphAr. GraphAr, on the other hand, can be considered as a new storage format with unique features tailored for LPGs. It can be leveraged by these works to further extend the utility and capabilities of data lakes.
Graph file formats. Certain formats are designed for graph (Brandes et al., 2002; Himsolt, 2010; gex, 2024; Batagelj and Mrvar, 2001) and RDF (Resource Description Framework) data (rdf, 2024; Martínez-Prieto et al., 2012). However, their primary focus is to describe or exchange data in a standardized manner, e.g., utilizing XML, and are not optimized for storage and retrieval purposes. The lack of encoding, compression and push-down optimizations can lead to far inferior performance, making them less suitable for managing LPGs in data lakes.
Graph-related databases. Some databases (neo, 2024b; Deutsch et al., 2019; jan, 2024; Feng et al., 2023) are designed to store and manage graph data. There are also efforts focus on optimizing graph-related queries (Nguyen et al., 2015; Zhao and Yu, 2017; Mhedhbi and Salihoğlu, 2022; Raasveldt and Mühleisen, 2019). While they offer various graph-related features, they primarily focus on in-memory mutable data management, operating at a higher level compared to GraphAr. GraphAr, with its format compatible with the LPG model, can be utilized as an archival format for graph databases.
Operation pushdown. Some previous works (Li et al., 2023; Willhalm et al., 2013, 2009; Yang et al., 2021; Yu et al., 2020; AWS, 2024) aim to develop high-performance operators on storage formats of either column-oriented or row-oriented. These works and GraphAr share the same goal of improving pushdown operators and leveraging CPU instructions like SIMD and BMI. However, these works mainly focus on operations related to relational data, such as scan, select, and filter based on properties. In contrast, GraphAr specifically focuses on two graph-specific operations.
8. Conclusion